# BIBO stability of port-Hamiltonian systems

### v1

@@ 3.3. Equivalent diagonal system
    It is a well-know result that any pHS satisfying Assumption 3.11 is regular and $L^{2}$-well-posed [12, Thm. 13.2.2].
@@
    Using the same splitting as in $P_{1}^{D}$ we can write $\mathcal{H}^{D}=\begin{pmatrix}\mathcal{H}^{D}_{+}&0\\ 0&\mathcal{H}^{D}_{-}\end{pmatrix}.$ Then +$\mathcal{H}^{D}_{+}$ contains the positive and $-\mathcal{H}^{D}_{-}$ the negative eigenvalues of $P_{1}\mathcal{H}$ on the respective diagonals. Furthermore the matrix $S^{-1}(\xi)$ then has the form
@@
    with $e_{1}(\xi),\ldots,e_{m}(\xi)$ eigenvectors of $P_{1}\mathcal{H}(\xi)$ corresponding to the eigenvalues in +$\mathcal{H}^{D}_{+}(\xi)$ and $e_{m+1}(\xi),\ldots,e_{n}(\xi)$ those corresponding to the ones in $-\mathcal{H}^{D}_{-}(\xi)$.
We now consider the state transformation $\widetilde{x}=Sx$. Then the PDE in Equation (2.1) describing the port-Hamiltonian system becomes
@@
Hence we find that the diagonalised system is just $\textrm{pH}_{n}\left(P_{1}^{D},P_{0}^{D},\mathcal{H}_{D},\widetilde{W}_{B}^{D},\widetilde{W}_{C}^{D}\right)$. As the system resulting from the state transformation has the same input-output behaviour as the original system, they have in particular the same transfer function and thus one is BIBO stable if and only if the other one is.
@@
Furthermore we find that the transformed input matrix $\widetilde{W}_{B}^{D}$ of the diagonalised system always possesses a decomposition of the form (3.7).
Proposition 3.13.
The matrix $\widetilde{W}^{D}_{B}\begin{bmatrix}P^{D}_{+}\\ P^{D}_{-}\end{bmatrix}$ is invertible and hence there exist matrices $K,M\in\mathbb{K}^{n\times n}$ with $K$ invertible such that
@@
where $\lambda_{i}(\xi)$ are the entries of ±$\mathcal{H}^{D}_{\pm}(\xi)$ corresponding to the eigenvectors $e_{i}(\xi)$ of $P_{1}\mathcal{H}(\xi)$.
@@
But by Theorem 1.5 in [11] this space is all of $\mathbb{K}^{n}$ as the operator $A$ generates a $C_{0}$-semigroup by Assumption 2.1. Hence $\widetilde{W}^{D}_{B}\begin{bmatrix}P^{D}_{+}\\ P^{D}_{-}\end{bmatrix}$ is invertible and existence of the decomposition follows from Proposition 3.3. ∎
Remark 3.14.
@@
    Using the explicit expressions from Proposition 3.3, we find that the matrix $M$ from Proposition 3.13 is given by
@@
    Note that the proof of Proposition 3.13 shows that the existence of the decomposition of $\widetilde{W}^{D}_{B}$ is in fact equivalent to Assumption 2.1.
@@ 3.4. The case $P^{D}_{0}=0$
We now want to restrict ourselves to a particular type of pHS and thus take the following assumption for the rest of this section.
Assumption 3.15.
$P_{0}^{D}=0$
Remark 3.16.
@@
    Note that Assumption 3.15 is satisfied in particular if the pHS has vanishing $P_{0}=0$ and a Hamiltonian $\mathcal{H}$ that is constant along $[a,b]$.
@@
    A concrete example on $[a,b]=[0,1]$ where this is the case despite non-constant $\mathcal{H}$ and $P_{0}\neq 0$ is given by $\textrm{pH}_{2}\left(\mathbbm{1}_{2},\left[\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right],\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ with
@@
which allows us to e.g. straightforwardly calculate the fundamental solution [^1] Note that in the following $\Psi_{\zeta}^{s}$, $\mathbf{V}(s)$, $\mathbf{U}(s)$, $M$, etc. will always refer to the objects calculated for the diagonalised system $\textrm{pH}_{n}\left(P_{1}^{D},P_{0}^{D},\mathcal{H}_{D},\widetilde{W}_{B}^{D},\widetilde{W}_{C}^{D}\right)$. to be
(3.10) $\Psi_{\zeta}^{s}=e^{s\left(P^{D}_{1}\right)^{-1}\int_{a}^{\zeta}\left(\mathcal{H}^{D}\right)^{-1}(\xi)\,\mathrm{d}\xi}=\begin{bmatrix}e^{s\int_{a}^{\zeta}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&e^{-s\int_{a}^{\zeta}\left(\mathcal{H}^{D}_{-}\right)^{-1}(\xi)\,\mathrm{d}\xi}\end{bmatrix}.$
@@
As a second step we can now also give explicit expressions for the matrices $\mathbf{V}(s)$, $\mathbf{V}(s)^{-1}$ and $\mathbf{U}(s)$ from Definition 3.8 appearing in the expression for the transfer function of the diagonalised system in Lemma 3.9,
(3.11) $\mathbf{V}(s)=\begin{bmatrix}e^{s\int_{a}^{b}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&\mathbbm{1}_{n-m}\end{bmatrix},\,\,\,\,\,\mathbf{V}(s)^{-1}=\begin{bmatrix}e^{-s\int_{a}^{b}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&\mathbbm{1}_{n-m}\end{bmatrix},$
(3.12) $\mathbf{U}(s)=\begin{bmatrix}e^{-s\int_{a}^{b}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&e^{-s\int_{a}^{b}\left(\mathcal{H}^{D}_{-}\right)^{-1}(\xi)\,\mathrm{d}\xi}\end{bmatrix}.$
@@
Lemma 3.17.
Let $\mathbf{U}(s)$ be given as in Equation (3.12).
@@
    The map $t\mapsto\|\mathbf{U}(t)\|_{\ell^{p}\rightarrow\ell^{p}}$ is strictly decreasing, $t\in\mathbb{R}$, and
    $\lim_{t\rightarrow\infty}\|\mathbf{U}(t)\|_{\ell^{p}\rightarrow\ell^{p}}=0$.
@@
whose form follows from Lemma 3.17(1). Then the properties of $\mathbf{U}(s)$ from that Lemma imply convergence of the Neumann series $\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}$ for certain $s$.
Lemma 3.18.
@@
    By Parts 5 to 7 of Lemma 3.17, there exists an $\alpha_{0}\geq 0$ such that $\|M\mathbf{U}(s)\|_{\ell^{2}\rightarrow\ell^{2}}\leq\|M\|_{\ell^{2}\rightarrow\ell^{2}}\|\mathbf{U}(s)\|_{\ell^{2}\rightarrow\ell^{2}}<1$ for $s\in\mathbb{C}_{\alpha_{0}}$. Hence the Neumann series $\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}$ converges for these $s$.
@@
Using Lemma 3.18 and Lemma 3.9 we then immediately have the following results.
Corollary 3.19.
@@
    (3.13) $\mathbf{G}(s)=\mathbf{Z}(s)\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}\,K^{-1}.$
@@
    If $\|M\|_{\ell^{2}\rightarrow\ell^{2}}\leq 1$ then (3.13) holds for $s\in\mathbb{C}_{0}$.
@@
    If $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$ then (3.13) holds for $s\in\overline{\mathbb{C}_{0}}$.
By Lemma 3.10(2) we thus have
$\mathcal{L}^{-1}\left\{\mathbf{G}\right\}(t)=\mathcal{L}^{-1}\left\{\mathbf{Z}\right\}(t)\ast\sum_{k=0}^{\infty}\mathcal{L}^{-1}\left\{M\mathbf{U}\right\}(t)^{\ast k}K^{-1}.$
@@ 3.5. BIBO stability of the diagonalised system
We now strive to derive sufficient conditions for BIBO stability using the condition from Part 3 of Lemma 3.10. By a slight abuse of notation aimed at conciseness, let in the following $\left\|F_{p,q}(s)\right\|_{\mathcal{M}}:=\left\|\mathcal{L}^{-1}\left\{F_{p,q}\right\}\right\|_{\mathcal{M}}$ denote the total variation of the inverse Laplace transform of the $(p,q)$-th entry of $F$.
With this notation, by Part 3 of Lemma 3.10 it is thus sufficient for BIBO stability to show that for all $1\leq p,q\leq n$ we have
@@
(3.14) $\sum_{k=0}^{\infty}\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,q}\right\|_{\mathcal{M}}<\infty.$
Lemma 3.21.
@@
and hence $\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,q}\right\|_{\mathcal{M}}=\sum_{i=1}^{m}\left|\alpha^{(k)}_{pq,i}\right|<\infty$ with $1\leq p,q\leq n$ always has bounded total variation. For sufficient conditions for BIBO stability it thus remains to guarantee that the sums in Equation (3.14) are all finite. We do this using a number of different additional assumptions on $M$.
Theorem 3.22.
Let $\textrm{pH}_{n}\left(P^{D}_{1},0,\mathcal{H}^{D},\widetilde{W}^{D}_{B},\widetilde{W}^{D}_{C}\right)$ be a pHS in diagonalised form satisfying Assumption 3.11. Then if any of the following conditions is satisfied, Equation (3.14) holds and hence the system is BIBO stable.
@@
    $\left(\left(M\mathbf{U}(s)\right)^{k+1}\right)_{p,q}$ $=\begin{bmatrix}\sum_{i}\alpha^{(k)}_{p1,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i}\alpha^{(k)}_{p2,i}e^{-s\widetilde{\lambda}_{i}}&\cdots&\sum_{i}\alpha^{(k)}_{pn,i}e^{-s\widetilde{\lambda}_{i}}\end{bmatrix}\begin{bmatrix}M_{1q}e^{-s\lambda_{q}}\\ M_{2q}e^{-s\lambda_{q}}\\ \vdots\\ M_{nq}e^{-s\lambda_{q}}\end{bmatrix}$
@@
    $\left(\left(M\mathbf{U}(s)\right)^{k+1}\right)_{p,q}$ $=\begin{bmatrix}\sum_{i}\alpha^{(k)}_{p1,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i}\alpha^{(k)}_{p2,i}e^{-s\widetilde{\lambda}_{i}}&\cdots&\sum_{i}\alpha^{(k)}_{pn,i}e^{-s\widetilde{\lambda}_{i}}\end{bmatrix}\begin{bmatrix}M_{1q}e^{-s\lambda_{q}}\\ M_{2q}e^{-s\lambda_{q}}\\ \vdots\\ M_{nq}e^{-s\lambda_{q}}\end{bmatrix}$
@@
    $\left\|\left(\left(M\mathbf{U}(s)\right)^{k+1}\right)_{p,q}\right\|_{\mathcal{M}}\leq\sum_{r}\sum_{i}\left|\alpha^{(k)}_{pr,i}M_{rq}\right|=\sum_{r}\left|M_{rq}\right|\sum_{i}\left|\alpha^{(k)}_{pr,i}\right|$
@@
    $\hskip 15.0pt=\begin{bmatrix}\sum_{i}\alpha^{(\ell k_{0})}_{p1,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i}\alpha^{(\ell k_{0})}_{p2,i}e^{-s\widetilde{\lambda}_{i}}&\cdots&\sum_{i}\alpha^{(\ell k_{0})}_{pn,i}e^{-s\widetilde{\lambda}_{i}}\end{bmatrix}\begin{bmatrix}\sum_{j}\alpha^{(k_{0})}_{1q,j}e^{-s\widetilde{\widetilde{\lambda}}_{j}}\\ \sum_{j}\alpha^{(k_{0})}_{2q,j}e^{-s\widetilde{\widetilde{\lambda}}_{j}}\\ \vdots\\ \sum_{j}\alpha^{(k_{0})}_{nq,j}e^{-s\widetilde{\widetilde{\lambda}}_{j}}\end{bmatrix}$
@@
    $\left\|\left(\left(M\mathbf{U}(s)\right)^{(\ell+1)k_{0}}\right)_{p,q}\right\|_{\mathcal{M}}$
    $\qquad\leq\sum_{r}\sum_{i}\sum_{j}\left|\alpha^{(\ell k_{0})}_{pr,i}\alpha^{(k_{0})}_{rq,j}\right|=\sum_{r}\sum_{i}\left|\alpha^{(\ell k_{0})}_{pr,i}\right|\left\|\left(\left(M\mathbf{U}(s)\right)^{k_{0}}\right)_{r,q}\right\|_{\mathcal{M}}.$
@@
It is easy to see that the norm condition from part (1) of Theorem 3.22 is only a sufficient but not a necessary condition.
Example 3.23.
Consider $P_{1}=\mathcal{H}=\mathbbm{1}_{2}$ and $M=\begin{bmatrix}\frac{3}{4}&\frac{1}{2}\\ 0&\frac{1}{2}\end{bmatrix}$. Then $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$ but $\|M\|_{\ell^{\infty}\rightarrow\ell^{\infty}}=\frac{5}{4}>1$. Explicit calculations show
$\left(M\mathbf{U}(s)\right)^{k}=\begin{bmatrix}\left(\frac{3}{4}\right)^{k}e^{-ks}&2^{1-k}\left(\left(\frac{3}{2}\right)^{k}-1\right)e^{-ks}\\ 0&\left(\frac{1}{2}\right)^{k}e^{-ks}\end{bmatrix}$
@@
and hence $\mathcal{L}^{-1}\left\{\left(M\mathbf{U}\right)^{k}\right\}(t)=\begin{bmatrix}\left(\frac{3}{4}\right)^{k}\delta(t-k)&2^{1-k}\left(\left(\frac{3}{2}\right)^{k}-1\right)\delta(t-k)\\ 0&\left(\frac{1}{2}\right)^{k}\delta(t-k)\end{bmatrix}.$
As $\sum_{k=0}^{\infty}\left(\frac{3}{4}\right)^{k}<\infty$, $\sum_{k=0}^{\infty}\left(\frac{1}{2}\right)^{k}<\infty$ and $\sum_{k=0}^{\infty}2^{1-k}\left(\left(\frac{3}{2}\right)^{k}-1\right)<\infty$, we can conclude that $\mathcal{L}^{-1}\left\{\mathbf{G}\right\}$ is a measure of bounded total variation and hence the system is BIBO stable.
@@
Furthermore, the following example provides a case where neither the sufficient condition from part (1) nor the one from part (2) of Theorem 3.22 are satisfied, but where part (3) can be applied.
Example 3.24.
Consider $\textrm{pH}_{2}\left(\mathbbm{1}_{2},0,2\mathbbm{1}_{2},\left[\begin{smallmatrix}1&0&\frac{1}{2}&\frac{1}{2}\\ 0&1&\frac{1}{2}&-\frac{1}{2}\end{smallmatrix}\right],\left[\begin{smallmatrix}0&0&1&0\\ 0&0&0&1\end{smallmatrix}\right]\right)$. Then we find

### v2

@@ 3.3. Equivalent diagonal system
    It is a well-know result that any hBCS satisfying Assumption 18 is regular and $L^{2}$-well-posed [21, Thm. 13.2.2].
@@
    Using the same splitting as in $P_{1}^{D}$ we can write $\mathcal{H}^{D}=\begin{pmatrix}\mathcal{H}^{D}_{+}&0\\ 0&\mathcal{H}^{D}_{-}\end{pmatrix}$. Then $\mathcal{H}^{D}_{+}$ contains the positive and $\mathcal{H}^{D}_{-}$ the negative eigenvalues of $P_{1}\mathcal{H}$ on the respective diagonals. Furthermore the matrix $S^{-1}(\xi)$ then has the form
@@
    with $e_{1}(\xi),\ldots,e_{m}(\xi)$ eigenvectors of $P_{1}\mathcal{H}(\xi)$ corresponding to the eigenvalues in $\mathcal{H}^{D}_{+}(\xi)$ and with $e_{m+1}(\xi),\ldots,e_{n}(\xi)$ those corresponding to the ones in $\mathcal{H}^{D}_{-}(\xi)$.
We now consider the state transformation $\widetilde{x}=Sx$. Then the PDE in Equation (1) describing the boundary control system becomes
@@
Hence we find that the diagonalised system is just $\Sigma_{n}(P_{1}^{D},P_{0}^{D},\mathcal{H}_{D},\widetilde{W}_{B}^{D},\widetilde{W}_{C}^{D})$. As the system resulting from the state transformation has the same input-output behaviour as the original system, they have in particular the same transfer function and thus one is BIBO stable if and only if the other one is.
@@
Furthermore we find that the transformed input matrix $\widetilde{W}_{B}^{D}$ of the diagonalised system always possesses a decomposition of the form (11).
Proposition 20.
Let $\Sigma_{n}(P_{1}^{D},P_{0}^{D},\mathcal{H}_{D},\widetilde{W}_{B}^{D},\widetilde{W}_{C}^{D})$ be the diagonalised version of a hBCS satisfying Assumptions 1 and 18. Then the matrix $\widetilde{W}^{D}_{B}\begin{bmatrix}P^{D}_{+}\\ P^{D}_{-}\end{bmatrix}$ is invertible and hence there exist matrices $K,M\in\mathbb{K}^{n\times n}$ with $K$ invertible such that
@@
where $\lambda_{i}(\xi)$ are the entries of $\mathcal{H}^{D}_{\pm}(\xi)$ corresponding to the eigenvectors $e_{i}(\xi)$ of $P_{1}\mathcal{H}(\xi)$.
@@
But by Theorem 1.5 in [18] this space is all of $\mathbb{K}^{n}$ as the operator AA generates a $C_{0}$-semigroup by Assumption 1(5). Hence $\widetilde{W}^{D}_{B}\begin{bmatrix}P^{D}_{+}\\ P^{D}_{-}\end{bmatrix}$ is invertible and existence of the decomposition follows from Proposition 10. ∎
Remark 21.
@@
    Using the explicit expressions from Proposition 10, we find that the matrix $M$ from Proposition 20 is given by
@@
    Note that the proof of Proposition 20 shows that the existence of the decomposition of $\widetilde{W}^{D}_{B}$ is in fact equivalent to Assumption 1(5).
@@ 3.4. The case $P^{D}_{0}=0$
We now want to restrict ourselves to a particular type of hyperbolic boundary control system and thus take the following assumption for the rest of this section.
Assumption 22.
For any $\xi\in[a,b]$ we have $P_{0}^{D}(\xi)=0$
Remark 23.
@@
    Note that Assumption 22 is satisfied in particular if the system has vanishing $P_{0}=0$ and a Hamiltonian $\mathcal{H}$ that is constant along $[a,b]$.
@@
    A concrete example on $[a,b]=[0,1]$ where this is the case despite non-constant $\mathcal{H}$ and $P_{0}\neq 0$ is given by $\Sigma_{2}\left(\mathbbm{1}_{2},\left[\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right],\mathcal{H},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ with
@@
which allows us to e.g. straightforwardly calculate the fundamental solution [^1] Note that in the following $\Psi_{\zeta}^{s}$, $\mathbf{V}(s)$, $\mathbf{U}(s)$, $M$, etc. will always refer to the objects calculated for the diagonalised system $\Sigma_{n}(P_{1}^{D},P_{0}^{D},\mathcal{H}_{D},\widetilde{W}_{B}^{D},\widetilde{W}_{C}^{D})$. to be
(14) $\Psi_{\zeta}^{s}=e^{s\left(P^{D}_{1}\right)^{-1}\int_{a}^{\zeta}\left(\mathcal{H}^{D}\right)^{-1}(\xi)\,\mathrm{d}\xi}=\begin{bmatrix}e^{s\int_{a}^{\zeta}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&e^{-s\int_{a}^{\zeta}\left(\mathcal{H}^{D}_{-}\right)^{-1}(\xi)\,\mathrm{d}\xi}\end{bmatrix}$.
@@
As a second step we can now also give explicit expressions for the matrices $\mathbf{V}(s)$, $\mathbf{V}(s)^{-1}$ and $\mathbf{U}(s)$ from Definition 15 appearing in the expression for the transfer function of the diagonalised system in Lemma 16,
(15) $\mathbf{V}(s)=\begin{bmatrix}e^{s\int_{a}^{b}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&\mathbbm{1}_{n-m}\end{bmatrix},\,\,\,\,\,\mathbf{V}(s)^{-1}=\begin{bmatrix}e^{-s\int_{a}^{b}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&\mathbbm{1}_{n-m}\end{bmatrix}$,
(16) $\mathbf{U}(s)=\begin{bmatrix}e^{-s\int_{a}^{b}\left(\mathcal{H}^{D}_{+}\right)^{-1}(\xi)\,\mathrm{d}\xi}&0\\ 0&e^{-s\int_{a}^{b}\left(\mathcal{H}^{D}_{-}\right)^{-1}(\xi)\,\mathrm{d}\xi}\end{bmatrix}$.
@@
Lemma 24.
Let $\mathbf{U}(s)$ be given as in Equation (16).
@@
    The map $t\mapsto\|\mathbf{U}(t)\|_{\ell^{p}\rightarrow\ell^{p}}$ is strictly decreasing, $t\in\mathbb{R}$, and $\lim_{t\rightarrow\infty}\|\mathbf{U}(t)\|_{\ell^{p}\rightarrow\ell^{p}}=0$.
@@
whose form follows from Lemma 24(1). Then the properties of $\mathbf{U}(s)$ from that Lemma imply convergence of the Neumann series $\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}$ for certain $s$.
Lemma 25.
Let $\mathbf{U}(s)$ be given as in Equation (16) and $M\in\mathbb{K}^{n\times n}$.
@@
    By Parts 5 to 7 of Lemma 24, there exists an $\alpha_{0}\geq 0$ such that $\|M\mathbf{U}(s)\|_{\ell^{2}\rightarrow\ell^{2}}\leq\|M\|_{\ell^{2}\rightarrow\ell^{2}}\|\mathbf{U}(s)\|_{\ell^{2}\rightarrow\ell^{2}}<1$ for $s\in\mathbb{C}_{\alpha_{0}}$. Hence the Neumann series $\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}$ converges for these $s$.
@@
Using Lemma 25 and Lemma 16 we then immediately have the following results.
Corollary 26.
Consider a hBCS satisfying Assumptions 1, 18 and 22.
@@
    (17) $\mathbf{G}(s)=\mathbf{Z}(s)\sum_{k=0}^{\infty}(M\mathbf{U}(s))^{k}\,K^{-1}.$
@@
    If $\|M\|_{\ell^{2}\rightarrow\ell^{2}}\leq 1$ then (17) holds for $s\in\mathbb{C}_{0}$.
@@
    If $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$ then (17) holds for $s\in\overline{\mathbb{C}_{0}}$.
By Lemma 17(2) we thus have
(18) $\mathcal{L}^{-1}\left\{\mathbf{G}\right\}(t)=\mathcal{L}^{-1}\left\{\mathbf{Z}\right\}(t)\ast\sum_{k=0}^{\infty}\mathcal{L}^{-1}\left\{M\mathbf{U}\right\}(t)^{\ast k}K^{-1}.$
@@ 3.5. BIBO stability of the diagonalised system
We now strive to derive sufficient conditions for BIBO stability using the condition from Part 3 of Lemma 17. By a slight abuse of notation aimed at conciseness, let in the following $\left\|F_{p,q}(s)\right\|_{\mathcal{M}}:=\left\|\mathcal{L}^{-1}\left\{F_{p,q}\right\}\right\|_{\mathcal{M}}$ denote the total variation of the inverse Laplace transform of the $(p,q)$-th entry of $F$.
With this notation, by Part 3 of Lemma 17 it is thus sufficient for BIBO stability to show that for all $1\leq p,q\leq n$ we have
@@
(19) $\sum_{k=0}^{\infty}\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,q}\right\|_{\mathcal{M}}<\infty.$
Lemma 28.
@@
[…]
Theorem 29.
Let $\Sigma_{n}\left(P^{D}_{1},0,\mathcal{H}^{D},\widetilde{W}^{D}_{B},\widetilde{W}^{D}_{C}\right)$ be the diagonalised form of a system satisfying Assumptions 1, 18 and 22. Then if any of the following conditions is satisfied, Equation (19) holds and hence the system is BIBO stable.
@@
    $\left(\left(M\mathbf{U}(s)\right)^{k+1}\right)_{p,q}$ $=\begin{bmatrix}\sum_{i}\alpha^{(k)}_{p1,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i}\alpha^{(k)}_{p2,i}e^{-s\widetilde{\lambda}_{i}}&\hskip-4.0pt\cdots\hskip-4.0pt&\sum_{i}\alpha^{(k)}_{pn,i}e^{-s\widetilde{\lambda}_{i}}\end{bmatrix}\begin{bmatrix}M_{1q}e^{-s\lambda_{q}}\\ M_{2q}e^{-s\lambda_{q}}\\ \vdots\\ M_{nq}e^{-s\lambda_{q}}\end{bmatrix}$
@@
    $\left(\left(M\mathbf{U}(s)\right)^{k+1}\right)_{p,q}$ $=\begin{bmatrix}\sum_{i}\alpha^{(k)}_{p1,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i}\alpha^{(k)}_{p2,i}e^{-s\widetilde{\lambda}_{i}}&\hskip-4.0pt\cdots\hskip-4.0pt&\sum_{i}\alpha^{(k)}_{pn,i}e^{-s\widetilde{\lambda}_{i}}\end{bmatrix}\begin{bmatrix}M_{1q}e^{-s\lambda_{q}}\\ M_{2q}e^{-s\lambda_{q}}\\ \vdots\\ M_{nq}e^{-s\lambda_{q}}\end{bmatrix}$
@@
[…]
@@
    $\hskip 15.0pt=\begin{bmatrix}\sum_{i}\alpha^{(\ell k_{0})}_{p1,i}e^{-s\widetilde{\lambda}_{i}}&\sum_{i}\alpha^{(\ell k_{0})}_{p2,i}e^{-s\widetilde{\lambda}_{i}}&\hskip-4.0pt\cdots\hskip-4.0pt&\sum_{i}\alpha^{(\ell k_{0})}_{pn,i}e^{-s\widetilde{\lambda}_{i}}\end{bmatrix}\begin{bmatrix}\sum_{j}\alpha^{(k_{0})}_{1q,j}e^{-s\widetilde{\widetilde{\lambda}}_{j}}\\ \sum_{j}\alpha^{(k_{0})}_{2q,j}e^{-s\widetilde{\widetilde{\lambda}}_{j}}\\ \vdots\\ \sum_{j}\alpha^{(k_{0})}_{nq,j}e^{-s\widetilde{\widetilde{\lambda}}_{j}}\end{bmatrix}$
@@
[…]
    $=\sum_{r}\sum_{i}\left|\alpha^{(\ell k_{0})}_{pr,i}\right|\left\|\left(\left(M\mathbf{U}(s)\right)^{k_{0}}\right)_{r,q}\right\|_{\mathcal{M}}.$
@@
Example 30.
  1. (1)
    As an example for a system where BIBO stability can be shown using the sufficient conditions from Theorem 29, take the following impedance passive port-Hamiltonian system from [21, Ex. 9.2.1] of a vibrating string with a damper affixed at one end (Figure 1).
    Figure 1. Vibrating string damped at the boundary
    Let the control inputs be given by the force $u_{1}$ applied at the right-hand side and by the velocity $u_{2}$ at the left-hand side. As observation outputs we choose the velocity $y_{1}$ at right-hand side and the force acting on the left-hand side. Then we can write this system as a pHS on the interval $[0,1]$ with
    $P_{1}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\quad P_{0}=0,\quad\mathcal{H}=\begin{bmatrix}\rho^{-1}&0\\ 0&T\end{bmatrix}$,
    $\widetilde{W}_{B}=\begin{bmatrix}k&1&0&0\\ 0&0&1&0\end{bmatrix},\quad\widetilde{W}_{C}=\begin{bmatrix}1&0&0&0\\ 0&0&0&1\end{bmatrix}$,
    where $\rho$ and $T$ are the mass density and strain of the string, and $k$ is the damping coefficient of the damper.
    Calculating the component of the transfer function from $u_{1}$ to $y_{1}$, we find it to be $\mathbf{G}_{11}(s)=\frac{1}{k+\sqrt{\rho T}\coth(\sqrt{\frac{\rho}{T}}s)}$. As no simple expression for the inverse Laplace transform of this function exists, we are again barred from applying [30, Thm. 3.4] directly and thus indeed have to resort to the sufficient conditions derived here.
    As this pHS satisfies Assumption 18, we can diagonalise it and find $M=\left[\begin{smallmatrix}0&\frac{k-\sqrt{\rho T}}{k+\sqrt{\rho T}}\\ 1&0\end{smallmatrix}\right]$. Now as $\|M\|_{\ell^{\infty}\rightarrow\ell^{\infty}}=\max(1,\frac{k-\sqrt{\rho T}}{k+\sqrt{\rho T}})=1\nless 1$, we cannot apply Theorem 29(1). However, we see that $\sum_{k=0}^{\infty}\left(M_{|\cdot|}\right)^{k}=\frac{1}{2\sqrt{\rho T}}\left[\begin{smallmatrix}k+\sqrt{\rho T}&k-\sqrt{\rho T}\\ k+\sqrt{\rho T}&k+\sqrt{\rho T}\end{smallmatrix}\right]$ converges and thus the system is BIBO stabe by Theorem 29(2).
    Using Equation (18) we can then even determine the inverse Laplace transform of $\mathbf{G}_{11}$ explicitly as
@@
    $\mathcal{L}^{-1}\left\{\mathbf{G}_{11}\right\}=\frac{1}{k+\sqrt{\rho T}}\delta(\cdot)-\frac{2\sqrt{\rho T}}{k^{2}-\rho T}\sum_{n=1}^{\infty}\left(\frac{k-\sqrt{\rho T}}{k+\sqrt{\rho T}}\right)^{n}\delta\left(\cdot-2n\sqrt{\frac{\rho}{T}}\right).$
    Note that this example further shows that the norm condition from part (1) of Theorem 29 is only a sufficient but not a necessary condition.
  2. (2)
@@
    There are systems for which neither the sufficient condition from part (1) nor the one from part (2) of Theorem 29 are satisfied, but where part (3) can be applied.